OpenStudy (anonymous):

factor the following sum of two cubes: m^3+64n^3

OpenStudy (callisto):

Use the identity : \(a^3 + b^3 = (a+b)(a^2-ab+b^2)\) Can you first identify what are a and b respectively?

OpenStudy (anonymous):

yeah, i know the format. i go it down to: (mn)^3+4^3

OpenStudy (callisto):

Nope.... 64 = x^3 What is x?

OpenStudy (anonymous):

4?

OpenStudy (callisto):

Yup. So, you get \[m^3+64n^3\]\[=m^3+(4n)^3\]You can apply the identity right away.

OpenStudy (anonymous):

is it: (m+4n)(m^2-4mn+4n^2)

OpenStudy (callisto):

Check the term "4n^2" again.

OpenStudy (callisto):

It should be(4n)^2, that is...

OpenStudy (anonymous):

eh, i got it wrong. It was (m+4)(m^2-4mn+16n^2)